Calculus Problem.
Prove that F(x) = 4x^3+x has no relative max or min without graphing.

Answer by stanbon(75887)

Prove that F(x) = 4x^3+x has no relative max or min without graphing.
----
F'(x) = 12x^2+1
-----
Solve: 12x^2+1 = 0
12x^2 = -1
x^2 = -1/12
------
Solutions for that are not Real Numbers.
Therefore F(x) has no relative max or min.
Cheers,
Stan H.
-----